OpenStudy (anonymous):

find zeros and state the multiplicity of zeros: y=x(x-1)^4

OpenStudy (anonymous):

when y=0 x(x-1)^4 = 0 for the equation to be 0, either x has to equal zero as 0*(-1)^4 = 0 or (x-1) has to equal 0 So, either x = 0 or x= 1 but x would = 1 4 times due to the fact that (x-1) is ^4

OpenStudy (anonymous):

if that makes sense?

OpenStudy (anonymous):

that makes sense. thanks

OpenStudy (anonymous):

No Problem :)